For a small manufacturing firm, the unit cost C(x) in dollars of producing x units per day is given by C(x)= x^2-120x+4000 how many items should be produced per day to minimize the unit cost, and what is the minimum unit cost?

Answer by stanbon(75887)

For a small manufacturing firm, the unit cost C(x) in dollars of producing x units per day is given by C(x)= x^2-120x+4000 how many items should be produced per day to minimize the unit cost, and what is the minimum unit cost?
----
minimum occurs at x = -b/(2a) = 120/(2*1) = 60 units per day
----
Minimum Cost = 60^2-120*60+4000 = $400
------------
Cheers,
Stan H.
-------------
----